# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v1

@@ 2.3.2 Physical Property-embedded Machine Learning
Physical properties usually include symmetry and conservation of physical systems. By building a neural network that inherently respects a given physical property, we thus make conservation of the associated quantity more likely and consequently the model’s prediction more physically accurate.
Conservation. Many real-world systems are governed by conservation laws related to mass, energy, momentum, charge, or particle counts, which are often expressed through continuity equations. For example, in hydrology, it is the amount of water [167]. However, standard deep learning approaches (such as CNNs and LSTM) struggle at conserving quantities across layers or timesteps. Thus, mass-conserving LSTM (MC-LSTM) [168] adheres to these conservation laws by extending the inductive bias of LSTM to model the redistribution of those stored quantities. MC-LSTMs [168] modify the recurrence in traditional LSTM to guarantee the conservation of the mass input. The key idea is to use the memory cells from LSTMs as mass accumulators, or mass storage. In addition, an antisymmetrical continuous convolutional layer [169] in a hierarchical network is presented by enforcing the conservation of momentum with a hard constraint. The parameterization of deep neural networks [170] is designed to satisfy the continuity equation, a fundamental conservation law, by building divergence-free neural networks. An implicit neural network layer [171] that incorporates the conservation of mass is proposed by using implicit differentiation. Furthermore, a systematic way of enforcing nonlinear analytic constraints (energy conservation and mass conservation) in neural networks via constraints in the architecture [172] is also proposed. The concept of “conversion layers” is introduced in the architecture that transforms nonlinearly constrained mappings into linearly-constrained mappings within NNs without overly degrading performance.
Symmetry has long been implicitly used in DL to design networks with known invariances and equivariances. Let $f:X\rightarrow Y$ be a function and $G$ be a group. Assume $G$ acts on $X$ and $Y$. If $f(gx)=gf(x)$ for all $x\in X$ and $g\in G$, the function $f$ is $G$-equivariant. If $f(gx)=f(x)$ for all $x\in X$ and $g\in G$, the function $f$ is G-invariant.
Convolutional neural networks enabled breakthroughs in computer vision by leveraging translation equivariance. Similarly, recurrent neural networks, graph neural networks, and capsule networks all impose symmetries [12]. There is a deep connection between symmetries and physics. For example, Noether’s law gives a correspondence between conserved quantities and groups of symmetries [173]. In addition, the symmetries of fluid dynamics [174] are exploited to design equivariant networks. The Navier-Stokes equations and heat equations are invariant under the following transformations:
@@
First, an end-to-end ML-based deep potential–smooth edition (DeepPot-SE) [177] is constructed. DeepPot-SE is extensive and continuously differentiable, scales linearly with system size, and preserves all the natural symmetries of the system. Translational, rotational, and permutational symmetries are preserved by symmetry preserving functions.
Second, there have been a number of GNN-based works, called geometrically equivariant graph neural networks [178], that leverage symmetry as an inductive bias in learning to simulate. These models are designed such that their outputs will rotate/translate/reflect in the same way as the inputs, hence retaining the symmetry. For example, EGNN [179] models the interaction with the invariant distance as input, computed as an inner product of the relative positions, and GMN [180] generalizes to a multi-channel version by taking into consideration a stack of multiple vectors. Subequivariant GNN [181] relaxes equivariance to subequivariance by considering external fields like gravity, where the universal approximation ability holds theoretically.
Third, an integral neural operator architecture [176] is designed to learn physical models with fundamental conservation laws automatically guaranteed. The translation- and rotation-invariant neural operator is developed by replacing the frame-dependent position information with its invariant counterpart in the kernel space; consequently it abides by the conservation laws of linear and angular momentums.
Causality. A fundamental pursuit in physics is to identify causal relationships. Integrating causality into ML promises to advance the comprehension of physical processes and bolster the robustness of ML models [182]. For example, causality-DeepONet [183] implements the physical causality in the DeepONet structure. Specifically, the causality-DeepONet is applied to learn the operator representing the response of a building due to earthquake ground accelerations. However, many questions of causality-embedded ML remain open, such as using causality to improve deep learning models and understand the system’s response under interventions, disentangling complex and multiple treatments, and designing the environment to control the given physics. In addition, a causal training algorithm  [184] is introduced to restore physical causality during the training of a PINNs model by appropriately re-weighting the PDE residual loss at each iteration of gradient descent. This simple modification provides PINNs with the capability to address complex problems, such as the incompressible Navier-Stokes equations in the turbulent regime.
@@ 2.3.3 Physical Condition-embedded Machine Learning
There are many physical conditions in the physical system, such as boundary conditions and initial conditions in PDEs, boundaries of land use and land cover in the physical environment, and so on. Violating these conditions can lead to unstable models and non-physical solutions. Thus, it is critical for ML to satisfy these constraints in order to capture the underlying physics accurately and provide reliable models for rigorous research and engineering design. Specifically, for PDEs, three types of boundary conditions (BCs) that are widely used to model physical phenomena are (1) Dirichlet: fixed-value at the boundary; (2) Neumann: fixed-derivative at the boundary; and (3) periodic: equal values at the boundary. In order to embed these BCs into NNs, a boundary enforcing operator network [185] is proposed that enables the BC satisfaction of neural operators by making structural changes to the operator kernel. The physics-embedded neural network [186] is developed by considering BCs in GNN.
@@ 2.4 Physics-aware Hybrid learning
PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models, with ML models. As shown in Fig. 4, according to the hybrid way, hybrid learning can be divided into serial, parallel, and complex ways of approaching learning.
@@ 2.4.1 Neural–numerical Hybrid Learning
Serial Hybrid has broad potential advantages for physical models and data-driven models. For example, NNs are used to handle data coming from irregular domains. A family of neural ODEs [187, 188] changes the traditionally discretized neuron layer depths into continuous equivalents such that the derivative of the hidden state can be parameterized using a neural network. The output of the network is then computed using a black box differential equation solver, making Neural ODEs an efficient combination of neural nets and numerical solvers. In addition, the advantages of physics solvers (such as their scale-invariant properties) can be imparted into NNs. A novel hybrid training approach [189] that combines a scale-invariant physics solver with ML techniques is developed for solving inverse problems, such as parameter estimation and optimal control. Gradients of NNs are replaced by the update computed from a higher-order solver that can encode the scale-invariant property of the physics.
Parallel Hybrid methods are designed to develop fast, robust, and reliable solvers for solving PDEs. Physics-driven models are usually used as a supervised signal to train data-driven NNs. For example, a hybrid, iterative, numerical, and transferable solver [190] called HINTS is proposed for differential equations. HINTS combines standard relaxation methods (such as Jacobi, Gauss-Seidel, conjugate gradient, multigrid methods, and their more advanced variants) and DeepONet. Compared to standard numerical solvers, HINTS is capable of providing faster solutions for a wide class of differential equations, while preserving accuracy close to machine zero. In addition, a hybrid graph neural network [191] that combines a traditional graph convolutional network with an embedded differentiable fluid dynamics simulator is developed. The hybrid graph neural network both generalizes well to new situations and benefits from the substantial speedup of neural network CFD predictions, while also substantially outperforming the coarse CFD simulation alone. A fast, differentiable hybrid approach [192] combines a 2D direct numerical simulation for the deformable solid structure and a physics-constrained neural network surrogate to capture the hydrodynamic effects of the fluid.
@@ 2.4.2 Neural–hydrology Hybrid Learning
The hydrological model, a standard representation of physical processes, serves as an input-output model simulating the evolution and dynamics of water storage, fluxes, and physical properties in the Earth’s surface and subsurface [2]. Neural-hydrology hybrid learning utilizes ML methods to harness and integrate a broad variety of hydrological models into a final process-based prediction product. Recently, neural–hydrology hybrid learning is receiving growing attention due to interpretability in hydrological models, a better appreciation of ML, and the availability of computing resources and methods [30]. Here we list some recent neural–hydrology hybrid methods for hydrological variable prediction, such as streamflow and runoff.
@@
Serial Hybrid combines the hydrological models and data-driven models sequentially. Serial hybrid includes a conceptual hydrological model that seeks to combine the strengths of data-driven and conceptual models, driven by dynamic predictions. For example, a hybrid monthly streamflow forecasting approach [194] is explored, where simulated soil moisture from the GR4J conceptual rainfall-runoff model is used to represent initial catchment conditions in a Bayesian NN model. In this work, the hybrid model performed better than either the GR4J model or the Bayesian neural network alone. In addition, to reduce the computational complexity of parallel hybridization of the conceptual rainfall-runoff models and ML techniques, NNs and support vector regression are embedded into a monthly lumped conceptual rainfall-runoff model [195]. Mohammadi et al. [196] use two process-driven conceptual rainfall-runoff models (HBV: Hydrologiska Byrans Vattenbalansavdelning, and NRECA: Non Recorded Catchment Areas) to provide inputs to support vector machines (SVMs) and an adaptive neuro-fuzzy inference system (ANFIS), in order to build seven variants of hybrid models. Results showed that AI-based hybrid models generally led to more accurate streamflow estimates compared with HBV and NRECA. In addition, three hybridization approaches that combine the simulations of the HBV model with LSTM [197] are developed for semi-arid regions. The results show distinct improvements in the three types of hybrid models when compared with the HBV model and standalone LSTM model. A number of studies have coupled conceptual models and data-driven models with serial hybrid methods.
@@
Parallel Hybrid. In a parallel hybrid structure, the data-driven and dynamic models are combined in parallel. Parallel hybrid models can employ a data-driven model to combine hydrological predictions in parallel. For example, a combination of conceptual and ML techniques [198] is proposed for enhancing the accuracy of streamflow simulation in a snow-covered basin. An end-to-end hybrid modeling approach [199] is presented to learn and predict spatial-temporal variations of observed and unobserved hydrological variables globally. The model combines a dynamic neural network and a conceptual water balance model, constrained by the water cycle observational products of evapotranspiration, runoff, snow-water equivalent, and terrestrial water storage variations. The model reproduces observed water cycle variations very well and the emergent relations of runoff-generating processes are qualitatively consistent with our understanding. Furthermore, the hybrid hydrological model [200] is used for global hydrological modeling that exploits the data adaptivity of neural networks for representing uncertain processes within a model structure based on physical principles (e.g., mass conservation). The hybrid model is trained simultaneously with a multi-task learning approach. It can conclude that the hybrid model provides a new data-driven perspective on modeling the global hydrological cycle and physical responses with machine-learned parameters that is consistent with and complementary to existing global modeling frameworks. In addition, a parallel hybrid approach can be employed to correct model biases [201].
@@
Complex Hybrid. In a complex structure, data-driven models can be integrated into the input space, output space, and interior of hydrological models in a complex hybrid method, according to different application requirements. Specifically, a general physics-AI approach [9] is proposed to improve AI geoscientific awareness, wherein temporal dynamic geoscientific models are included as a special process-wrapped recurrent neural network (denoted as P-RNN). In addition, the common NN layers without physics address processes unrepresented by the P-RNN layer. The experiments of the physics-AI approach demonstrate that the complex hybrid model has enhanced prediction accuracy, robust transferability, and good intelligence for inferring unobserved processes. Furthermore, a differentiable programming framework [202] is designed to parameterize, enhance, or replace the process-based model’s modules in a simple hydrologic model, HBV, for predictions of hydrologic variables (streamflow). Specifically, an LSTM outputs the physical parameters for a process-based model. In process-based models, certain parts of the model can be replaced by neural networks and the structure can be updated. The framework is trained in an end-to-end fashion and there are no intermediate supervising data (or labels) for the outputs of the neural networks. The main loss function is calculated between the process-based model output and observations.
@@ 2.5 Summary of Physics-aware Machine Learning
Physical Data-guided Machine Learning (PDgML)
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data.
Classical physical data-guided neural networks (spatio-temporal modeling, dynamic system modeling), deep operator networks, physics-discovery neural networks
[Benefits] Only need simulations from physics model, fast inference, the discovery of the dynamic behavior;
@@
Physics-informed Machine Learning (PiML)
PiML is a widely used approach to incorporating physical constraints that can be trained from additional information obtained by enforcing the physical laws (for example, designing loss functions or regularization).
Physics-informed neural networks (PINNs), physics-informed deep neural operators (PINOs)
[Benefits] Incomplete models and imperfect data, strong generalization in small data regime, understanding deep learning by the optimization process, tackling high dimensionality; uncertainty quantification, data-physics-driven parameters discovery;
@@
Physics-embedded Machine Learning (PeML)
PeML is achieved by embedding physics in the model frameworks or modules.
Physical equation-embedded machine learning (PEeML), physical property-embedded machine learning (PPeML), physical condition-embedded machine learning (PCeML)
[Benefits] Interpretability of architecture and learning process, generalization;
@@
Physics-aware Hybrid Learning (PaHL)
PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models with ML models. Depending on the hybrid approach, hybrid learning can be categorized as serial, parallel, or complex.
Neural–numerical hybrid learning, neural–hydrology hybrid learning
[Benefits] Can leverage both the flexibility of data-driven models and the interpretability and generalizability of physics-driven models, speed, and operational convenience;

### v4

@@ 2.3.2 Physical Property-embedded Machine Learning
Physical properties frequently encompass symmetries and conservation laws inherent in physical systems. Constructing a neural network that inherently respects a specific physical property enhances the likelihood of conserving relevant quantities, thereby improving the model’s predictions in terms of physical accuracy.
Conservation. Many real-world systems adhere to conservation laws related to mass, energy, momentum, or particle number, typically expressed through continuity equations. For example, in hydrology, it is the amount of water [167]. However, standard deep learning methods, such as CNNs and LSTMs, encounter challenges in maintaining conservation across layers or time steps. To address this issue, the mass-conserving LSTM (MC-LSTM) [168] extends the inductive bias of the LSTM to uphold these conservation laws, ensuring that mass input is conserved through modifications to the recurrent structure of the traditional LSTM. The main concept is to utilize memory cells from LSTMs as mass accumulators or storage units. In addition, an antisymmetrical continuous convolutional layer [169] in a hierarchical network is presented by enforcing the conservation of momentum with a hard constraint. The parameterization of deep neural networks [170] is tailored to meet the continuity equation, ensuring the adherence to fundamental conservation laws by creating divergence-free NNs. An implicit neural network layer [171] that incorporates the conservation of mass is proposed by using implicit differentiation. Furthermore, a structured approach to enforce nonlinear analytic constraints such as energy and mass conservation within NNs [172] is proposed. The concept of “conversion layers” is introduced within the architecture to transform nonlinearly constrained mappings into linearly constrained ones within NNs, achieving this without significantly compromising performance.
Symmetry has long been implicitly utilized in DL for designing networks with known invariances and equivariances. Let $f:X\rightarrow Y$ be a function and $G$ be a group. Assume $G$ acts on $X$ and $Y$. If $f(gx)=gf(x)$ for all $x\in X$ and $g\in G$, the function $f$ is $G$-equivariant. If $f(gx)=f(x)$ for all $x\in X$ and $g\in G$, the function $f$ is G-invariant.
Convolutional neural networks have revolutionized computer vision by utilizing translation equivariance. Similarly, symmetries are enforced in recurrent neural networks, graph neural networks, and capsule networks [12]. There exists a profound connection between symmetry and physics. For instance, Noether’s theorem establishes a relationship between conserved quantities and symmetry groups [173]. Moreover, symmetries in fluid dynamics [174] are employed in the design of equivariant networks. The Navier-Stokes equations and the heat equation exhibit invariance under the following transformations:
@@
First, an end-to-end ML-based deep potential–smooth edition (DeepPot-SE) [177] is developed, which is extensive, continuously differentiable, and scales linearly with system size while preserving all natural symmetries. The symmetry-preserving functions maintain translational, rotational, and permutational symmetries.
Second, there have been a number of GNN-based works, called geometrically equivariant graph neural networks [178], which leverage symmetry as an inductive bias in learning simulations. These models ensure that their outputs rotate, translate, or reflect in the same manner as their inputs, preserving symmetry.For example, EGNN [179] models interactions using invariant distances, calculated via inner products of relative positions, while GMN [180] extends to a multi-channel version with a stack of vectors. Subequivariant GNN [181] relaxes the condition of full equivariance to subequivariance, accommodating external fields like gravity while retaining the theoretical ability for universal approximation.
Thirdly, an integral neural operator architecture [176] is designed to automatically learn physical models with fundamental conservation laws guaranteed. This translation- and rotation-invariant neural operator replaces frame-dependent position information with invariant counterparts in kernel space, ensuring conservation laws of linear and angular momenta are upheld.
Causality. A fundamental pursuit in physics is to identify causal relationships. Integrating causality into ML promises to advance the comprehension of physical processes and bolster the robustness of ML models [182]. For example, causality-DeepONet [183] implements the physical causality in the DeepONet structure. Specifically, the causality-DeepONet is employed to learn operators that capture the response of buildings to earthquake ground acceleration. However, many challenges in causal embedding in ML remain unresolved. These include leveraging causality to enhance deep learning models, understanding system responses under interventions, disentangling complex and multiple processes, and designing environments to control physical phenomena. Additionally, a causal training algorithm [184] has been introduced to restore physical causality during PINNs model training by appropriately reweighting the PDE residual loss in each iteration of gradient descent. This straightforward adjustment enables PINNs to address complex problems such as the incompressible Navier-Stokes equations in turbulent regimes.
@@ 2.3.3 Physical Condition-embedded Machine Learning
There are many physical conditions in the physical system, such as boundary conditions and initial conditions in PDEs, boundaries of land use and land cover in the physical environment, and so on. Failing to satisfy these conditions can result in unstable models and non-physical solutions. Therefore, it is essential for ML models to adhere to these constraints to accurately capture underlying physical principles and provide reliable models for scientific and engineering applications. Specifically, for PDEs, three types of boundary conditions (BCs) that are widely used to model physical phenomena are (1) Dirichlet: fixed-value at the boundary; (2) Neumann: fixed-derivative at the boundary; and (3) periodic: equal values at the boundary. To integrate boundary conditions into NNs, a boundary enforcing operator network [185] is introduced. This network ensures the satisfaction of boundary conditions by modifying the operator kernel’s structure. The physics-embedded neural network [186] is developed by considering BCs in GNN.
@@ 2.4 Physics-aware Hybrid learning
PaHL directly combines pure physics-based models, such as numerical methods, climate, hydrology, land, and Earth system models, with ML models. As shown in Fig. 4, according to the hybrid way, hybrid learning can be divided into serial, parallel, and complex ways of approaching learning.
@@ 2.4.1 Neural–numerical Hybrid Learning
Serial Hybrid has broad potential advantages for physical models and data-driven models. For example, NNs are used to handle data coming from irregular domains. A family of neural ODEs [187, 188] transforms traditionally discretized neuron layer depths into continuous equivalents by parameterizing the derivative of the hidden state with a neural network. The network’s output is then computed using a black-box differential equation solver, effectively merging NNs with numerical solvers for efficiency. In addition, the advantages of physics solvers (such as their scale-invariant properties) can be imparted into NNs. A novel hybrid training approach [189] that combines a scale-invariant physics solver with ML techniques is developed for solving inverse problems, like parameter estimation and optimal control. Gradients of NNs are replaced by the update computed from a higher-order solver that can encode the scale-invariant property of the physics.
Parallel Hybrid methods are designed to develop fast, robust, and reliable solvers for solving PDEs. Physics-driven models are usually used as a supervised signal to train data-driven NNs. For example, a solver named HINTS (Hybrid, Iterative, Numerical, and Transferable Solver) [190] is proposed for differential equations. HINTS combines traditional relaxation techniques (such as Jacobi, Gauss-Seidel, conjugate gradient, multigrid methods, and their variants) with DeepONet. This approach offers faster solutions for a wide range of differential equations while maintaining machine-level accuracy. Additionally, a hybrid graph neural network [191] that merges a traditional graph convolutional network with an integrated differentiable fluid dynamics simulator has been developed. This hybrid model generalizes well to new scenarios and benefits from the significant speedup of neural network-based CFD predictions, significantly outperforming standalone coarse CFD simulations. A fast, differentiable hybrid approach [192] integrates a 2D direct numerical simulation for deformable solid structures with a physics-constrained neural network surrogate to capture fluid hydrodynamic effects.
@@ 2.4.2 Neural–hydrology Hybrid Learning
The hydrological model is a standard representation of physical processes, serving as an input-output model used to simulate the evolution and dynamics of surface and groundwater storage, fluxes, and physical properties of the Earth [2]. Neural-hydrology hybrid learning integrates various hydrological models using machine learning methods to produce a final process-based prediction product. This approach has garnered increased attention recently due to its ability to enhance the interpretability of hydrological models, improve understanding of machine learning techniques, and leverage advances in computational resources and methods [30]. Here we list some recent neural–hydrology hybrid methods for hydrological variable prediction, such as streamflow and runoff.
@@
Serial Hybrid integrates hydrological models and data-driven models in a sequential manner. For instance, the hybrid monthly runoff forecasting method [194] utilizes simulated soil moisture from the GR4J conceptual rainfall-runoff model to represent initial catchment conditions within a Bayesian neural network framework. This serial hybrid model demonstrates superior performance compared to both the standalone GR4J model and the Bayesian neural network. Additionally, to mitigate the computational complexity associated with parallel hybridization of conceptual rainfall-runoff models and ML techniques, Okkan et al. [195] embeds NNs and support vector regression into a monthly centralized conceptual rainfall-runoff model. Furthermore, Mohammadi et al. [196] use two process-driven conceptual rainfall-runoff models—HBV (Hydrologiska Byrans Vattenbalansavdelning) and NRECA (Non Recorded Catchment Areas)—to provide inputs for support vector machines (SVMs) and an adaptive neuro-fuzzy inference system (ANFIS), resulting in seven hybrid model variants. The results indicate that AI-based hybrid models generally deliver more accurate streamflow estimates compared to the HBV and NRECA models alone. Moreover, three hybrid methods integrating HBV model simulations with LSTM [197] are developed for semi-arid regions, demonstrating significant improvements over both the HBV and standalone LSTM models. Numerous studies illustrate the efficacy of serial hybrid methods in combining conceptual hydrological models with data-driven models.
@@
Parallel Hybrid. In a parallel hybrid architecture, data-driven models and dynamic models are integrated concurrently. This approach can employ a data-driven model to combine hydrological predictions in parallel. For instance, a hybrid approach combining conceptual and ML techniques has been proposed to enhance the accuracy of runoff simulations in snow-covered basins [198]. An end-to-end hybrid modeling approach is introduced to learn and predict the spatiotemporal variations of observed and unobserved hydrological variables globally [199]. This model integrates a dynamic neural network with a conceptual water balance model, constrained by water cycle observational products such as evapotranspiration, runoff, snow water equivalent, and changes in terrestrial water storage. The model accurately reproduces observed water cycle variations, and the relationships of runoff generation processes are well aligned with established understanding. Furthermore, a hybrid hydrological model [200] is used for global hydrological modeling, utilizing neural networks’ adaptability to represent uncertain processes within a framework grounded in physical principles, such as mass conservation. This hybrid model is simultaneously trained using a multi-task learning approach, indicating that it provides a novel data-driven perspective for modeling the global hydrological cycle and physical responses through machine-learned parameters. This method aligns with existing global modeling frameworks. Moreover, parallel hybrid methods have been shown to correct model biases [201].
@@
Complex Hybrid. In a complex structure, data-driven models can be integrated into the input space, output space, and interior of hydrological models in a complex hybrid method, according to different application requirements. Specifically, a general physics-AI approach [9] is proposed to improve AI geoscientific awareness, wherein temporal dynamic geoscientific models are included as a special process-wrapped recurrent neural network (denoted as P-RNN). In addition, the common NN layers that lack physics-based processes address those unrepresented by the P-RNN layer. Experiments with the physics-AI approach show that this complex hybrid model improves prediction accuracy, robustness in transferability, and the ability to infer unobserved processes.Additionally, a differentiable programming framework [202] is designed to parameterize, enhance, or replace the process-based model’s modules in a simple hydrologic model, HBV, for predicting hydrologic variables such as streamflow. Specifically, the LSTM outputs the physical parameters of the process-based model. Within the process-based model, certain components can be substituted with NNs and the structure is updatable. The framework is trained end-to-end, without intermediate supervisory data or labels. The primary loss function is computed between the output of the process-based model and the observed data.
@@ 2.5 Summary of Physics-aware Machine Learning
Physical Data-guided Machine Learning (PDgML) PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data. Classical physical data-guided neural networks (spatio-temporal modeling, dynamic system modeling), deep operator networks, physics-discovery neural networks [Benefits] Only need simulations from physics model, fast inference, the discovery of the dynamic behavior;
@@
Physics-informed Machine Learning (PiML) PiML is a widely used approach to incorporating physical constraints that can be trained from additional information obtained by enforcing the physical laws (for example, designing loss functions or regularization). Physics-informed neural networks (PINNs), physics-informed deep neural operators (PINOs) [Benefits] Incomplete models and imperfect data, strong generalization in the limited data, understanding deep learning by the optimization process, tackling high dimensionality; uncertainty quantification, data-physics-driven parameters discovery;
@@
Physics-embedded Machine Learning (PeML) PeML is achieved by embedding physics in the model frameworks or modules. Physical equation-embedded machine learning (PEeML), physical property-embedded machine learning (PPeML), physical condition-embedded machine learning (PCeML) [Benefits] Interpretability of architecture and learning process, generalization;
@@
Physics-aware Hybrid Learning (PaHL) PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models with ML models. Depending on the hybrid approach, hybrid learning can be categorized as serial, parallel, or complex. Neural–numerical hybrid learning, neural–hydrology hybrid learning [Benefits] Can leverage both the flexibility of data-driven models and the interpretability and generalizability of physics-driven models, speed, and operational convenience;